OpenStudy (anonymous):

Solve the following system of equations: 3x + 6y = 9 x + 2y = 3 A. No solution B. Infinitely many solutions C. (3, 6) D. (1, 2)

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

i dont know if its no solution or infinitely many.......

OpenStudy (uri):

3x + 6y = 9 x + 2y = 3(multiply with -3) 3x + 6y = 9 -3x-2y=-9 ===== everything will be cancelled...so no solution.